OpenStudy (anonymous):

A 50.0 g projectile is launched with a horizontal velocity of 647 m/s from a 4.65 kg launcher moving in the same direction at 2.00 m/s. What is the launchers velocity after the launch?

OpenStudy (anonymous):

Mp Vp1 + ML VL 1 = Mp Vp2 + ML VL2 50*2+ 4.65*2 = 50*647 = 2*VL2 trickiest part is to realize taht is the projectile was traveling int he launcher, they both started out with velocity of 2 m/s. VL2 = -16120.35 m/s ridiculous number, but imagine, i have a launcher and it is tiny. and it is carrying MASSIVE weight to launch. and then it launches it at a ginormous velocity. i will expect the launcher to go flying backwards. hope that makes sense!